Using the following year-end information for Calvin’s Clothing, calculate the current ratio and acid-test ratio for the business:

Cash $ 51,680
Short-term investments 8,000
Accounts receivable 46,000
Inventory 150,000
Prepaid expenses 8,520
Accounts payable 102,000
Other current payables 30,100

rev: 07_10_2018_QC_CS-131052

  • 1.00 and 1.25

  • .50 and 1.25

  • 2.50 and 1.25

  • 2.00 and 0.80

  • 1.00 and 0.80
Answer #1

Current assets=Cash+Short-term investments+AR+Inventory+Prepaid expenses

=(51680+8000+46000+150,000+8520)=$264200

Current liabilities=Accounts payable+Other current payables

=(102000+30100)=$132100

Current ratio=Current assets/Current liabilities

=(264200/132100)=2

Acid-test ratio=(Current assets-Inventory-Prepaid expenses)/Current liabilities

=(264200-8520-150,000)/132100)=0.8

Hence the correct option is 2.00 and 0.80